# Localizing Factual Inconsistencies in Attributable Text Generation

### v1

@@ 4.2 Advantages of QASemConsistency
We measure inter annotator agreement using Fleiss Kappa on all annotated QAs. We achieve a $\kappa$ = 0.72 for XSUM and $\kappa=0.79$ for FActScore, indicating a substantial agreement. For comparison, previous factuality benchmarks report lower agreement: Pagnoni et al. (2021) report $\kappa=0.58$ at the sentence-level and  Cao and Wang (2021) reports $\kappa=0.35$ at the token-level. Annotating factual consistency is a challenging and sometimes subjective task (Falke et al., 2019). We believe that the high agreement is due to the QASemConsistency’s decomposition into atomic semantic relations, where annotators need to assess a single assertion at a time. In addition, by collecting factual consistency annotation for each predicate-argument relation, QASemConsistency ensures that the human raters will assess all details.
QASemConsistency vs. Token-Level Annotations.
@@
We compare the localized annotations of supported QAs in the CLIFF’s subset of our QASemConsistency dataset to the original token-level annotations in CLIFF. Specifically, we analyze the cases where summaries were annotated as fully supported in CLIFF but not in our annotations and vice-versa. Notably, we observe that all ten summaries where all QAs were annotated as supported in our dataset were also annotated as fully supported in CLIFF. Conversely, we identified 10 out of 74 summaries marked as fully supported in CLIFF whereas our annotators found unsupported QAs. Upon analyzing these cases, we discover that the majority of the summaries (7 out of 10) were indeed not fully supported by the reference text, as reflected by the unsupported QAs. For instance, the summary sentence “The US has suspended its participation in talks with Russia to try to broker a cessation of hostilities in Syria, the State Department says.” is not fully supported by the reference text “The United States is suspending its participation in bilateral channels with Russia that were established to sustain the cessation of hostilities” because the reference text mentions that the suspension of the U.S participation is “to sustain the cessation” and not “to broker a cessation”. This fine-grained factual inconsistency was indicated by several QAs marked as not supported, such as “Why has someone suspended something? To try to broker a cessation of hostilities in Syria” about the predicate “suspended”. Two other summaries could be interpreted as either supported or not supported, with our annotators showing localized disagreement on these cases. One summary was mistakenly annotated as not supported by our annotators. This analysis confirms that having human annotators verify each predicate-argument relation is not only beneficial for localizing factual inconsistencies, but also effectively helps annotators to identify them.
4.3 Evaluation Benchmark
Dataset #Summaries #QAs IAA
XSUM 74 692 0.72
FActScore 36 935 0.79
Table 3: Statistics of our collected benchmark QASemConsistency.
Overall, we gathered a dataset of entailment annotations for 1627 QASem QAs (see Table 3 for more statistics). The ground truth label for each QA is determined by the majority vote among the annotators. This dataset allows evaluating the performance of automatic detectors of localized factual inconsistencies, as we do in §5, and is released (upon publication) to facilitate future research of inconsistency detection models.
@@ 5 Automatic Detection of Localized Factual Inconsistencies
In this section we describe several models that automatically localize factual inconsistencies according to our methodology. We decompose the generated text $\mathcal{Y}$ into a list of QA pairs $\{qa_{1},qa_{2},...,qa_{n}\}$, and define the likelihood for each $qa_{i}$ to be supported (entailed) by the reference text $\mathcal{X}$ as $s(\mathcal{X},qa_{i})\in[0,1]$. We use the QASem parser from  Section 3 to generate the QAs and conduct our experiments with different entailment classifiers.
@@ 5.1 Entailment Classifiers
We apply two recent off-the-shelf NLI models to predict whether the QASem QA $qa_{i}$ is entailed by the reference text. In these experiments, the premise is the reference text $\mathcal{X}$ and the hypothesis is the concatenation of the question $q_{i}$ and the answer $a_{i}$ in $qa_{i}$.
The first model is DeBERTa-NLI, an encoder-only model based on DeBERTA-V3-Large (He et al., 2023), which was trained over a mixture of challenging NLI datasets, including MNLI (Williams et al., 2018), ANLI (Nie et al., 2020), FEVER (Thorne et al., 2018), WNLI (Liu et al., 2022) and LingNLI (Parrish et al., 2021). [^7] https://huggingface.co/MoritzLaurer/DeBERTa-v3-large-mnli-fever-anli-ling-wanli This model includes a classification head that classifies the input to entailment, neutral and contradiction. We consider the probability of the entailment class as our score $s(\mathcal{X},qa_{i})$.
The second supervised classifier is from TRUE (Honovich et al., 2022), an encoder-decoder model based on T5-XXL (Raffel et al., 2020) and finetuned on diverse entailment datasets: SNLI (Bowman et al., 2015), MNLI (Williams et al., 2018), FEVER (Thorne et al., 2018), Scitail (Khot et al., 2018), PAWS (Zhang et al., 2019) and VitaminC (Schuster et al., 2021). [^8] https://huggingface.co/google/t5_xxl_true_nli_mixture The TRUE model, also known as AutoAIS (Bohnet et al., 2022), was widely used in previous work for automatically measuring factual consistency (Bohnet et al., 2022; Gao et al., 2023a; Roit et al., 2023; Gao et al., 2023b; Slobodkin et al., 2024). This model was trained to generate “1” if the hypothesis is entailed by the premise and “0” otherwise. Hence, $s(\mathcal{X},qa_{i})$ is set as the probability of generating “1”, given the likelihood score of “1” and “0”.
@@
We also use several LLMs to predict $s(\mathcal{X},qa_{i})$ using zero-shot prompting. [^9] We also experimented with few-shot demonstrations on a sample of the QASemConsistency dataset but these did not improve performance. We instruct the model to write “Yes” if the proposition corresponding to the QA is entailed by the reference text, and “No” otherwise. The prompt used in our experiments is presented in Appendix C. We apply this prompt to recent open-source LLMs including Llama 8B (v3.1) (Dubey et al., 2024), Gemma 7B (v1.1) (Team et al., 2024), Gemma 9B (v2) (Riviere et al., 2024), Mistral 7B (v0.2) (Jiang et al., 2023) and Mixtral 8x7B (v0.1) (Jiang et al., 2024). We do not use commercial models because our evaluation is based on the probability score $s(\mathcal{X},qa_{i})\in[0,1]$ of generating “Yes” or “No” (see §5.2). For all prompted models, we define the entailment score $s$ as the probability of generating “Yes” as the first generated token, based on the likelihood score of generating “Yes” and “No”.
@@ 5.2 Evaluation of Automatic Detection
We apply the above models to predict an entailment score $s(\mathcal{X},qa_{i})$ for all QAs in our collected QASemConsistency dataset and compare the prediction to the ground truth. We report two metrics to evaluate the performance of our automatic models in detecting localized factual inconsistencies. First, we follow standard evaluation practices (Honovich et al., 2022) and report the Area Under the Receiver Operating Characteristic Curve (ROC AUC), which plots the true positive rate against the false positive rate for different possible thresholds of $s(\mathcal{X},qa_{i})$. Second, since the models are trained or instructed to classify a QA as either supported or not supported, we also measure the performance of the hard prediction, considering a QA as supported if $s(\mathcal{X},qa_{i})\geq 0.5$. As our main goal is to localize unsupported QAs, which becomes more and more subtle in recent models, we propose to report the F1 score for identifying unsupported QAs. That is, we consider a true positive instance when both the ground truth and the model predict “not supported”.
@@ 5.3 Results
CLIFF FActScore
F1 AUC F1 AUC
Supervised NLI models
DeBERTa-NLI (304M) 37.2 67.9 47.0 75.9
TRUE (11B) 62.2 69.9 65.5 86.4
LLM prompting
Gemma v1.1 (7B) 52.4 72.7 71.3 84.1
Gemma v2 (9B) 53.7 78.3 76.6 90.2
Mistral v0.3 (7B) 43.9 71.7 67.8 84.1
Mixtral v0.1 (8x7B) 57.5 69.2 70.8 80.6
Llama v3.1 (7B) 47.6 74.3 72.4 88.9
Table 4: Results of automatic evaluators on our collected benchmark QASemConsistency.
Table 4 presents the results of our automatic evaluation. Notably, Gemma v2 (9B) achieves the highest ROC AUC scores, reaching 78.3 on CLIFF and 90.2 on FActScore. The TRUE model achieves the highest F1 score on CLIFF (62.2) and Gemma v2 (9B) achieves the highest F1 score on FActScore (76.6). These results suggest that even though these models were trained on standard entailment datasets, where the hypothesis is a single sentence, they can effectively generalize to the hypotheses in our setting, that is, QASem QAs that represent predicate-argument relations (e.g, “How someone died? from measles”). This ability is likely as such models were trained on a large amount of text while QASem QAs are also natural language expressions.
CLIFF FActScore
Prec. Rec. Prec. Rec.
Supervised NLI models
DeBERTa-NLI (304M) 73.1 25.8 73.7 36.4
TRUE (11B) 46.7 93.2 51.0 91.2
LLM prompting
Gemma v1.1 (7B) 69.7 42.0 71.7 70.9
Gemma v2 (9B) 83.0 39.7 88.5 67.5
Mistral v0.3 (7B) 74.2 31.2 75.9 70.9
Mixtral v0.1 (8x7B) 70.8 48.5 62.6 81.6
Llama v3.1 (8B) 81.8 33.6 84.5 63.3
Table 5: Precision and Recall scores for automatic evaluators on the QASemConsistency dataset. For most models, precision is substantially higher than recall.
Observing the F1 scores, the best F1 score on CLIFF is 62.2 and the best F1 score on FActScore is 76.6. In particular, as detailed in Table 5, precision tends to be significantly higher than recall, often with a substantial gap. For instance, Gemma 2 (9B) achieves a precision of 83.0 on CLIFF but a recall of only 39.7. This discrepancy suggests that LLMs may be over confident when generating a “Yes” response to indicate that a QA is supported. Hence, instead of relying solely on the model’s binary predictions, we suggest that users adjust the decision threshold in order to optimize performance.
@@ 6 Related Work
Starting with coarse-grained evaluation, SummEval (Fabbri et al., 2021) asks annotators to assign a 1-5 likert score to the entire summary, while some other works aim to produce a single score/label for the entire output (Yin et al., 2021; Rashkin et al., 2021; Honovich et al., 2022; Tang et al., 2023; Liu et al., 2023b). Several works evaluate each generated sentence separately with some (simple or sophisticated) form of aggregation (Falke et al., 2019; Kryscinski et al., 2020; Pagnoni et al., 2021; Utama et al., 2022; Tang et al., 2022; Laban et al., 2022; Gekhman et al., 2023; Mishra et al., 2024; Subbiah et al., 2024).
To achieve sub-sentence evaluation, a few recent works decompose each sentence into “claims”, “facts” or “propositions” (Table 1 Claim-level), whose support by the reference text is then assessed independently (Min et al., 2023; Krishna et al., 2023; Chen et al., 2023; Kamoi et al., 2023a; Wanner et al., 2024). Such decomposition has been also proposed for evaluating the summary content relevance, following the Pyramid paradigm (Nenkova and Passonneau, 2004; Liu et al., 2023c; Wan et al., 2024). However, these free text claims typically lack structure and systematicity. As a result, identifying unsupported claims does not effectively localize factual mistakes in the generated text. In fact, these claims are not atomic and encompass multiple semantic relations. For instance, the claim “The man died of measles” in Table 1 is not supported by the article, but it remains unclear whether the issue is that the man died for a reason other than measles, or that there is no man who died.
@@
Some other works ask human annotators to highlight inconsistent tokens or spans in the generated text (Maynez et al., 2020; Cao and Wang, 2021). This method often results in relatively poor inter-annotator agreement (e.g., 0.35 Fleiss Kappa in CLIFF), because span annotation is subjective (Mishra et al., 2024) and individual spans might serve multiple roles in the sentence, where some roles are in correct assertions and some are not. Indeed, the token-level evaluation in Table 1 includes the tokens “how he got the illness”, although this phrase also implies that the man got the illness, which is supported by the reference text. In that summary, the mistake is that the goal of the examination should be to determine the cause of death, rather than how the man contracted the illness. This unsupported fact cannot be captured with span highlighting while it can effectively identified using our QA “What didn’t something establish? how he got the illness” (Table 1). In the same line of work,  Laban et al. (2023) create SummEdits, a challenging benchmark with localized factual inconsistencies. However, unlike QASemConsistency, the mistakes are limited to a single token and are not naturally occurring.
Finally, Goyal and Durrett (2020) propose DAE, an automatic evaluation metric that operates at the level of semantic dependency arcs in a structured semantic dependency representation (Oepen et al., 2014) to localize inconsistent semantic relations in the generated text. Similarly, FactGraph (Ribeiro et al., 2022) represents both the source article and the summary with AMR (Banarescu et al., 2013) then model the factual inconsistencies at the edge level. However, these methods have only been applied automatically because obtaining human judgments at this level of granularity is challenging. Indeed, annotators would need to understand dependency or AMR labels and isolate the semantics of individual arcs within sentences, making manual evaluation difficult. Similarly to DAE and FactGraph, our QASemConsistency also operates at the level of individual semantic relations, while representing them with simple natural language question-answer pairs, enabling human evaluation as well and easing LLM processing. Finally, Cho et al. (2024) shown success at applying Neo-Davidsonian formal semantics to automatic text-to-image evaluation.

### v2

@@ 4.2 Dataset Properties
We evaluate inter-annotator agreement using Fleiss’ Kappa. We observed substantial agreement across all three tasks: $\kappa=0.72$ for CLIFF, $\kappa=0.79$ for FactScore, and $\kappa=0.67$ for Verifiability. For comparison, previous factuality benchmarks report lower agreement: Pagnoni et al. (2021) report $\kappa=0.58$ at the sentence-level and  Cao and Wang (2021) reports $\kappa=0.35$ at the token-level. Annotating factual consistency is a challenging and sometimes subjective task (Falke et al., 2019). We believe that the high agreement is due to the QASemConsistency’s decomposition into atomic semantic relations, where annotators need to assess a single assertion at a time. In addition, by collecting factual consistency annotation for each predicate-argument relation, QASemConsistency ensures that the human raters will assess all details.
4.3 Analysis
4.3.1 Overall Score
Figure 3: Visualization of the correlation between the difference in factual consistency scores (calculated using QASemConsistency and token-based approach for “Cliff”) and human side-by-side preferences.
Here, we aim to show that QASemConsistency not only enables the localization of factual inconsistency but also provides an overall score that reflects well the degree of inconsistency in the response. To achieve this, we collect side-by-side annotations of a source text $x$ paired with two different model-generated responses $y_{1}$ and $y_{2}$. Annotators were asked to compare the factual consistency of the two responses on a scale of 1 to 5, where 1 indicates that $y_{2}$ is much more consistent than $y_{1}$, 2 that $y_{2}$ is more consistent than $y_{2}$, 3 that they are almost equivalent and symmetrically 4 and 5 indicate advantage of $y_{1}$ over $y_{2}$. This pairwise comparative judgment (Thurstone, 1927; David, 1963), a common and intuitive approach for comparing model outputs, allows human annotators to reliably compare outputs on a shared scale (Chiang et al., 2024).
Given a scoring function $s(x,y)$ that assigns a factual consistency score to the output $y$ with respect to the reference text $x$, we expect the difference $d(y_{1},y_{2})=s(x,y_{1})-s(x,y_{2})$ to correlate with the side-by-side score. Indeed, if $y_{1}$ is much more consistent than $y_{2}$, we expect $d$ to be a large positive value, whereas if the two responses are nearly equivalent in consistency, $d$ should be near zero.
We collect these side-by-side annotations on the summaries from CLIFF (37 pairs) and the biographies from FactScore (16 pairs). [^7] We cannot do this analysis for verifiability because different models generate responses that point to different sources. As mentioned above (§3), we define $s_{QA}(x,y)$ as the percentage of supported QAs in $y$ and compute this score using human labels. The Spearman correlation between $d(y_{1},y_{2})$ and the side-by-side consistency score yields strong positive results. For the summaries, we obtain a high correlation of $\rho=0.73$ (p-value << 0.001), while the token-level evaluation in CLIFF ($\rho=0.11$) and QGQA ($\rho=0.2)$ show no correlation. Figure 3 plots the difference in factual consistency scores derived from QASemConsistency and from token-level annotations against human side-by-side preferences. For the biographies, we obtain a high correlation for all metrics ($\rho=0.71$ for QASemConsistency, $\rho=0.67$ for FactScore, $\rho=0.81$ for QGQA). However, due to the overlapping confidence intervals and the small sample size ($N=16$), we cannot conclusively determine that one metric is statistically superior to the others based solely on these results. These strong correlations demonstrate that QASemConsistency effectively reflects the degree of inconsistency in model-generated responses and can be reliably used for ranking models.
4.3.2 Qualitative Analysis
@@
We compare the localized annotations of supported QAs in the CLIFF’s subset of our dataset to the original token-level annotations in CLIFF. Specifically, we analyze the cases where summaries were annotated as fully supported in CLIFF but not in our annotations and vice-versa. Notably, we observe that all ten summaries where all QAs were annotated as supported in our dataset were also annotated as fully supported in CLIFF. Conversely, we identified 10 out of 74 summaries marked as fully supported in CLIFF whereas our annotators found unsupported QAs. Upon analyzing these cases, we discover that the majority of the summaries (7 out of 10) were indeed not fully supported by the reference text, as reflected by the unsupported QAs. For instance, the summary sentence “The US has suspended its participation in talks with Russia to try to broker a cessation of hostilities in Syria, the State Department says.” is not fully supported by the reference text “The United States is suspending its participation in bilateral channels with Russia that were established to sustain the cessation of hostilities” because the reference text mentions that the suspension of the U.S participation is “to sustain the cessation” and not “to broker a cessation”. This fine-grained factual inconsistency was indicated by several QAs marked as not supported, such as “Why has someone suspended something? To try to broker a cessation of hostilities in Syria” about the predicate “suspended”. Two other summaries could be interpreted as either supported or not supported, with our annotators showing localized disagreement on these cases. One summary was mistakenly annotated as not supported by our annotators. This analysis confirms that having human annotators verify each predicate-argument relation is not only beneficial for localizing factual inconsistencies, but also effectively helps annotators to identify them more accurately.
@@ 5 Automatic Detection of Localized Factual Inconsistencies
In this section we describe several models that automatically localize factual inconsistencies according to our methodology. We decompose the generated text $y$ into a list of QA pairs $\{qa_{1},qa_{2},...,qa_{n}\}$, and define the likelihood for each $qa_{i}$ to be supported (entailed) by the reference text $x$ as $s(x,qa_{i})\in[0,1]$. We use the QASem parser from  Section 3 to generate the QAs and conduct our experiments with different entailment classifiers.
@@ 5.1 Entailment Classifiers
CLIFF FActScore Generative Search
BAcc AUC BAcc AUC BAcc AUC
Supervised NLI models
TRUE (11B) 61.4 72.1 72.5 88.8 68.5 79.1
TrueTeacher (11B) 68.9 82.1 71.7 88.7 57.8 80.5
Minicheck (770M) 70.4 80.8 78.0 87.3 76.3 83.4
LLM prompting
Gemma 2 2B 65.9 73.7 74.6 81.9 70.8 80.2
Gemma 2 9B 70.0 82.5 78.0 88.1 77.1 87.1
Gemma 2 27B 72.3 85.2 75.2 90.4 71.6 85.5
Mistral Nemo (12.2B) 70.6 79.5 73.7 85.4 71.5 82.4
Llama 3.1 8B 65.8 79.0 76.5 89.4 68.5 83.7
GPT-4o 75.7 - 82.4 - 79.2 -
Table 4: Performance of automatic models on the test set of our collected benchmark. We cannot report AUC for GPT-4o because it does not directly output probability scores for “Yes” or “No” answers. The highest BAcc and AUC within each category (Supervised NLI model and LLM prompting) are shown in bold.
@@
We apply three recent off-the-shelf NLI models to predict whether the QASem QA $qa_{i}$ is entailed by the reference text. In these experiments, the premise is the reference text $x$ and the hypothesis is the concatenation of the question $q_{i}$ and the answer $a_{i}$ in $qa_{i}$.
The first supervised classifier is TRUE (Honovich et al., 2022), an encoder-decoder model based on T5-XXL (11B parameters) (Raffel et al., 2020) and finetuned on diverse entailment datasets: SNLI (Bowman et al., 2015), MNLI (Williams et al., 2018), FEVER (Thorne et al., 2018), Scitail (Khot et al., 2018), PAWS (Zhang et al., 2019) and VitaminC (Schuster et al., 2021). [^8] https://huggingface.co/google/t5_xxl_true_nli_mixture The TRUE model, also known as AutoAIS (Bohnet et al., 2022), was widely used in previous work for automatically measuring factual consistency (Bohnet et al., 2022; Gao et al., 2023a; Roit et al., 2023; Gao et al., 2023b; Slobodkin et al., 2024). This model was trained to generate “1” if the hypothesis is entailed by the premise and “0” otherwise. Hence, $s(x,qa_{i})$ is set as the probability of generating “1”, given the likelihood score of “1” and “0”. The second supervised model is TrueTeacher (Gekhman et al., 2023), a T5-XXL model finetuned on many model generated summaries annotated for factual consistency using an LLM. Similarly to TRUE, this model was trained to generated “1” if the hypothesis is entailed by the premise. The last supervised model is MiniCheck (Tang et al., 2024), a recent T5 Large (770M parameters) trained on synthetic data generated with GPT-4. This model outperforms all systems of comparable size and reaches GPT-4 accuracy on LLM-AggreFact.
@@
We also use several LLMs to predict $s(x,qa_{i})$ using few-shot prompting. We instruct the model to write “Yes” if the proposition corresponding to the QA is entailed by the reference text, and “No” otherwise. The prompt used in our experiments is presented in Appendix C. We first apply this prompt to recent open-source LLMs including Llama 8B (v3.1) (Dubey et al., 2024), Gemma (v2) 2B, 9B and 27B (Riviere et al., 2024) and Mistral Nemo Instruct [^9] https://mistral.ai/news/mistral-nemo. For all prompted models, we define the entailment score $s$ as the probability of generating “Yes” as the first generated token, based on the likelihood score of generating “Yes” and “No”. Finally, we also use GPT-4o to predict entailment for the QAs but evaluate only “hard” predictions (see Section 5.2) because it does not output distributions over the vocabulary.
@@ 5.2 Evaluation of Automatic Detection
We report two common metrics to evaluate the performance of automatic models to predict entailment of QAs. First, we follow standard evaluation practices (Honovich et al., 2022; Gekhman et al., 2023) and report the Area Under the Receiver Operating Characteristic Curve (ROC AUC), which plots the true positive rate against the false positive rate for different possible thresholds of $s(x,qa_{i})$. Second, since the models are trained or instructed to classify a QA as either supported or not supported, we also measure the performance of the hard prediction. Similarly to (Tang et al., 2024), we do not perform threshold tuning for each dataset and consider a QA as supported if $s(x,qa_{i})\geq 0.5$. Following previous work (Laban et al., 2022; Fabbri et al., 2022; Tang et al., 2023, 2024), we report the Balanced accuracy metric (BAcc): $BAcc=\frac{1}{2}(\frac{TP}{TP+FN}+\frac{TN}{TN+FP})$.
@@ 5.3 Results
Table 4 presents the results of our automatic models. Notably, GPT-4o achieves the highest balanced accuracy (BAcc) across all datasets (75.7 on CLIFF, 82.4 on FactScore and 79.2 on Verifiability). Open source LLMs and supervised NLI models achieve also decent performance, with Minicheck emerging as the top-performing supervised model. These results suggest that even though these models were trained on standard entailment datasets, where the hypothesis is a single sentence, they can effectively adapt to our setting, where hypotheses take the form of question-answer pairs representing predicate-argument relations (e.g, “How someone died? from measles”). This generalization ability likely stems from the models’ training on massive text corpora, coupled with the fact that QASem question-answer pairs are expressed in natural language, unlike traditional semantic role labeling (SRL) schemes like PropBank or FrameNet that rely on predefined and complex taxonomies. Finally, the consistently higher performance on the FactScore dataset can be explained by the nature of its biographical content. These texts contain a high frequency of copular sentences (e.g., “Roselyn Sanchez is an actress.”), which present a simpler verification challenge for the models. Indeed, 15% of the QAs in FactScore are copular constructions (e.g, “Who is an actress? Roselyn Sanchez”, for which GPT-4o achieves a BAcc of 0.89, compared to 0.81 for non-copula QAs.
Beyond standard metrics, we want to assess how well model’s overall scores align with human judgment, similar to our approach with human scores (see Section 4.3.1). We compute the Spearman correlation between the difference of the individual QA scores $d(y_{1},y_{2})=s_{QA}(x,y_{2})-s_{QA}(x,y_{1})$ and the side-by-side human preference Likert scale. Our experiments show that GPT-4o achieves a Spearman correlation of $\rho=0.54$ (p-value < 0.05) on CLIFF. [^10] The results on FactScore were not statistically significant due to the limited number of pairwise preferences ($n=16$).
While GPT-4o achieves a high performance across datasets, there is still a gap between automatic and human performance, leaving much room for improvement in future work.
5.4 Analysis
QA vs. affirmative.
To investigate the impact of the question-answer (QA) format on entailment performance, we convert the QA pairs into affirmative sentences (e.g, “Who ate something? John” to “John ate something”) using a small LLM (Gemma 2 2B). We then prompt Gemma 2 9B to determine whether these affirmative sentences are supported by the reference text. Surprisingly, this approach results in substantially lower balanced accuracy (BAcc) scores compared to the QA format: 63.7 (-6.3) on CLIFF, 74.7 (-3.3) on FActScore and 73.4 (-3.7) on Verifiability. We believe that the QA format provides more structured information than simple affirmative sentences because (1) it explicitly delineates the predicate (within the question) and the argument (within the answer), and (2) the question formulation (e.g., “who”, “where”, etc.) offers an additional layer of semantic information which can be valuable for deciding entailment.
@@ 6 Related Work
Starting with coarse-grained evaluation, SummEval (Fabbri et al., 2021) asks annotators to assign a 1-5 likert score to the entire summary, while some other works aim to produce a single score/label for the entire output (Yin et al., 2021; Rashkin et al., 2023; Honovich et al., 2022; Tang et al., 2023; Liu et al., 2023b; Gekhman et al., 2023). Several works evaluate each generated sentence separately with some (simple or sophisticated) form of aggregation (Falke et al., 2019; Kryscinski et al., 2020; Pagnoni et al., 2021; Utama et al., 2022; Tang et al., 2022; Laban et al., 2022; Mishra et al., 2024; Subbiah et al., 2024; Tang et al., 2024).
To achieve sub-sentence evaluation, a few recent works decompose each sentence into “claims”, “facts” or “propositions” (Table 1 Claim-level), whose support by the reference text is then assessed independently (Min et al., 2023; Krishna et al., 2023; Chen et al., 2023; Kamoi et al., 2023a; Wanner et al., 2024; Samir et al., 2024; Wei et al., 2024; Wan et al., 2024). However, these free text claims typically lack structure and systematicity. As a result, identifying unsupported claims does not effectively localize factual mistakes in the generated text. In fact, these claims are not atomic and encompass multiple semantic relations. For instance, the claim “The man died of measles” in Table 1 is not supported by the article, but it remains unclear whether the issue is that the man died for a reason other than measles, or that there is no man who died.
@@
Some other works ask human annotators to highlight inconsistent tokens or spans in the generated text (Maynez et al., 2020; Cao and Wang, 2021). This method often results in relatively poor inter-annotator agreement (e.g., 0.35 Fleiss Kappa in CLIFF), because span annotation is subjective (Mishra et al., 2024) and individual spans might serve multiple roles in the sentence, where some roles are in correct assertions and some are not. Indeed, the token-level evaluation in Table 1 includes the tokens “how he got the illness”, although this phrase also implies that the man got the illness, which is supported by the reference text. In that summary, the mistake is that the goal of the examination should be to determine the cause of death, rather than how the man contracted the illness. This unsupported fact cannot be captured with span highlighting while it can effectively identified using our QA “What didn’t something establish? how he got the illness” (Table 1). In the same line of work,  Laban et al. (2023) create SummEdits, a challenging benchmark with localized factual inconsistencies. However, unlike our benchmark, the mistakes are limited to a single token and are not naturally occurring.
Goyal and Durrett (2020) propose DAE, an automatic evaluation metric that operates at the level of semantic dependency arcs in a structured semantic dependency representation (Oepen et al., 2014) to localize inconsistent semantic relations in the generated text. Similarly, FactGraph (Ribeiro et al., 2022) represents both the source article and the summary with AMR (Banarescu et al., 2013) then model the factual inconsistencies at the edge level. However, these methods have only been applied automatically because obtaining human judgments at this level of granularity is challenging. Indeed, annotators would need to understand dependency or AMR labels and isolate the semantics of individual arcs within sentences, making manual evaluation difficult. Similarly to DAE and FactGraph, our QASemConsistency also operates at the level of individual semantic relations, while representing them with simple natural language question-answer pairs, enabling human evaluation as well and easing LLM processing. Cho et al. (2024) shown success at applying Neo-Davidsonian formal semantics to automatic text-to-image evaluation.
We note though that while QASemConsistency is a promising approach for both manual and automatic evaluation, it focuses solely on measuring factual consistency, a crucial aspect of attributable text generation. To provide a more comprehensive assessment, we suggest that future work use QASemConsistency in addition to targetted metrics for capturing other aspects, such as BERTScore (Zhang* et al., 2020) for relevance, BookScore (Chang et al., 2024) for coherence, or with LLM as a judge (Zheng et al., 2023).
Furthermore, it is worth noting that recent advancements in reference-based metrics for the relevance aspect, such as Pyramid (Nenkova and Passonneau, 2004), LitePyramid (Shapira et al., 2019) and RoSE (Liu et al., 2023c), highlight the growing interest in more granular evaluation of relevance. Since QASemConsistency decomposes text into fine-grained predicate-argument assertions, we conjecture that it could be adapted also for providing more fine-grained evaluation of relevance.